# Dissipation meets conformal interface: How the relaxation rate is suppressed

### v1

@@ V Suppression of relaxation for characterizing conformal interface
which is calculated to first order in perturbation theory. That is, (40) only depends on the interface parameter $\lambda$. This motives us to define the relaxation coefficient $c_{\text{relax}}$ as
@@
where $c_{\text{relax}}^{(1)}$ denotes the approximation of $c_{\text{relax}}$ up to the first order perturbation. It is expected that (42) holds for small dissipation strength $\gamma$ in the perturbative regime $\gamma\ll J$, where $J=1$ is the hopping strength in (6). Physically, $c_{\text{relax}}$ captures how the relaxation rate in dissipative systems is suppressed by the presence of a conformal interface, similar to how the energy and information transmissions are suppressed in closed systems. As shown in Fig. 1, we compare $c_{\text{relax}}$ with $c_{\text{LR}}$ and $c_{\text{eff}}$ in the critical free fermion chain. One can find they show distinct features from each other – this indicates that $c_{\text{relax}}$ is indeed a new quantity characterizing the conformal interface in a dissipative system.
@@
Figure 7: Numerical results for the rescaled Liouvillian gap $(2N)^{3}g(\lambda)$ and $(2N)^{3}g(\lambda)/c^{(1)}_{\text{relax}}=(2N)^{3}g(\lambda)/(1-\sqrt{1-\lambda^{2}})$, plotted as a function of the dissipation strength $\gamma$. (a) and (b): The local dissipation is applied at the boundary, i.e. at $n_{d}=1$. (c) and (d): The local dissipation is applied right next to the interface, i.e. at $n_{d}=N-1$. In (b), the rescaled Liouvillian gap collapses to a single curve of $\pi^{2}\gamma/(1+\gamma^{2})$, denoted by the dashed line. In (d), the weak-dissipation regime is highlighted, where $c_{\text{relax}}$ exhibits a linear dependence on $\gamma$ and all curves collapse to the same one in this regime. The total system size is chosen as $2002N=$ 200.
@@
As a short summary, by comparing with the exact numerical calculation, we have verified our analytical result in (42) in the weak dissipation regime. That is, $c_{\text{relax}}$ is independent of both the dissipation strength $\gamma$ and the dissipation location $n_{d}$, as long as the dissipation is weak ($\gamma\ll J$). Very interestingly, when the dissipation is at the boundary, our result in (42) holds for arbitrarily finite dissipation strengths $\gamma$. This interesting feature may deserve a future study.
VI Discussion and conclusion
@@ Appendix A Lattice calculation of $c_{\text{LR}}$
Figure 8: Energy scattering by a conformal interface after a local quantum quench. The sample plots here correspond to the conformal interfaces with total transmission ($\lambda=1$), partial transmission/partial reflection ($\lambda=0.6$), and total reflection ($\lambda=1$). Here $\lambda$ is introduced in (6) and (7). We join the two decoupled chains at $x=-50$, and the conformal interface is inserted at $x=0$. The two decoupled chains before joint together are defined on $[-400,-50]$ and $[-50,400]$, respectively.
Next, we will use (44) and (45) to extract $c_{\text{LR}}$ on a lattice model. The incident energy pulse in (43) will be generated by a local quantum quench. More concretely, we consider two decoupled free-fermion chains, one defined on sites from $-N$ to $-N+N_{1}$, and the other defined on sites from $-N+N_{1}+1$ to $N$.
The conformal interface is inserted locally at two sites 00 and $1$. That is, the Hamiltonians for the two decoupled chains are $H_{1}$ and $H_{2}$, where
@@
where $\{c_{i}^{\dagger},c_{j}\}=\delta_{ij}$ and $\{c_{i},c_{j}\}=\{c_{i}^{\dagger},c_{j}^{\dagger}\}=0$. $H_{2}$ has the same expression as (6), except that now the system is defined on sites from $-N+N_{1}+1$ to $N$ and the conformal defect is defined on sites 00 and $1$.
@@
where $H_{\text{int}}=-\frac{1}{2}c_{N_{1}}^{\dagger}c_{N_{1}+1}+h.c.$. As shown in Fig.8, the excitations are generated from $x=-N+N_{1}$, where we join the two chains. As time evolves, we have a right-moving pulse and a left-moving pulse. When the right-moving energy pulse arrives at the conformal interface (at $x=0$), it will split into two parts: one is transmitted and the other is reflected. The amount of energy transmitted/reflected depends on the details of the conformal interface, which is characterized by $\lambda$ in (7). As shown in Fig.8, when the conformal interface is totally transmissive ($\lambda=1$), the right-moving energy pulse propagates freely in space. If the conformal interface is partially transmissive ($10<\lambda<$ 1), the pulse is partially transmitted and partially reflected. If the interface is totally reflective ($\lambda=0$), then one can find that the energy pulse is totally reflected, which becomes left-moving.
@@ Appendix B Lattice calculation of $c_{\text{eff}}$
First, we consider two decoupled CFTs defined in $[-l,0]$ and $[0,l]$ and join them at their ends at $x=0$ through a conformal interface. As shown in [54], the time evolution of entanglement entropy for the subsystem $A=[0,l]$ is described by
@@ Appendix C Liouvillian gap for general locations
Figure 11: The Liouvillian gap $g(\lambda)$ as a function of the total system size $L=2N$, with local dissipation applied at (a) $n_{d}=N-10$ and (b) $n_{d}=\frac{N}{2}$. For both the cases with ($\lambda=0.8$) and without ($\lambda=1$) conformal interface, we observe good cubic scaling, suggesting that it universally holds for general choices of the dissipative site $n_{d}$. The total system size is chosen as $8002N=200,260,320,\dots,$ 800, and the dissipation strength is set to a small value of $\gamma=0.05$.
@@ Appendix D Dynamics of particle density
Figure 12: Time evolution of particle density $\langle c_{x}^{\dagger}c_{x}\rangle_{t}$ in the presence of local dissipation at a site $n_{d}$ in the left half chain. (a) The dissipation is applied at the left boundary $n_{d}=1$, and we examine the particle density at the same site $x=1$. (b) The dissipation is applied at at the left boundary $n_{d}=1$, and we examine the particle density at site $x=N+10=110$. (c) The dissipation is applied at the closest site to the interface $n_{d}=N-1$, and we examine the particle density at at the left boundary $x=1$. (d) The dissipation is applied at the closest site to the interface $n_{d}=N-1$, and we examine the particle density at site $x=N+10=110$. Here, the total system is of length $2002N=$ 200 and the dissipation strength is set to $\gamma_{-}=0.1$. Here the initial states are set to be the ground state of the interface Hamiltonian, therefore the initial values of the particle density varies with the interface parameter $\lambda$.
@@
[…]
@@
[…]
where $\textbf{H}_{\text{eff}}=\textbf{H}-\frac{i}{2}(\mathbf{\Gamma}^{+}+\mathbf{\Gamma}^{-})$ is the (kernel of the) effective non-hermitian Hamiltonian as defined in Eq. (31) in the main text (but including both gain and loss here), and $\mathbf{\Gamma}^{\pm}_{x,y}=\gamma_{\pm}\delta_{x,y}\delta_{x,n_{d}}$ describes the particle gain and loss at the same site $n_{d}$.

### v2

@@ VI Suppression of relaxation for characterizing conformal interface
which is calculated to first order in perturbation theory in $\gamma$ (and also for the finite value of $\gamma$ when the dissipation is located at the boundary of a long critical chain). That is, (54) only depends on the interface parameter $\lambda$. This motives us to define the relaxation coefficient $c_{\text{relax}}$ as
@@
where $c_{\text{relax}}^{(1)}$ denotes the approximation of $c_{\text{relax}}$ up to the first order perturbation. It is expected that (56) holds for small dissipation strength $\gamma$ in the perturbative regime $\gamma\ll J$, where $J=1$ is the hopping strength in (6). Moreover, when local dissipation is introduced precisely at one of the boundaries, we expect the result (56) to hold for any finite dissipation strength $\gamma$. Physically, $c_{\text{relax}}$ captures how the relaxation rate in dissipative systems is suppressed by the presence of a conformal interface, similar to how the energy and information transmissions are suppressed in closed systems. As shown in Fig. 1, we compare $c_{\text{relax}}$ with $c_{\text{LR}}$ and $c_{\text{eff}}$ in the critical free fermion chain. One can find they show distinct features from each other – this indicates that $c_{\text{relax}}$ is indeed a new quantity characterizing the conformal interface in a dissipative system.
@@
Figure 7: Numerical results for the rescaled Liouvillian gap $(2N)^{3}g(\lambda)$ and $(2N)^{3}g(\lambda)/c^{(1)}_{\text{relax}}=(2N)^{3}g(\lambda)/(1-\sqrt{1-\lambda^{2}})$, plotted as a function of the dissipation strength $\gamma$. (a) and (b): The local dissipation is applied at the boundary, i.e. at $n_{d}=1$. (c) and (d): The local dissipation is applied right next to the interface, i.e. at $n_{d}=N-1$. In (b), the rescaled Liouvillian gap collapses to a single curve of $\pi^{2}\gamma/(1+\gamma^{2})$, denoted by the dashed line. In (d), the weak-dissipation regime is highlighted, where $c_{\text{relax}}$ exhibits a linear dependence on $\gamma$ and all curves collapse to the same one in this regime. The total system size is chosen as $2N=200$.
@@
As a short summary, by comparing with the exact numerical calculation, we have verified our analytical result in (56) in the weak dissipation regime. That is, $c_{\text{relax}}$ is independent of both the dissipation strength $\gamma$ and the dissipation location $n_{d}$, as long as the dissipation is weak ($\gamma\ll J$). Very interestingly, when the dissipation is at the boundary, our result in (56) holds for arbitrarily finite dissipation strengths $\gamma$.
VII Discussion and conclusion
@@ Appendix A Lattice calculation of $c_{\text{LR}}$
Figure 8: Energy scattering by a conformal interface after a local quantum quench. The sample plots here correspond to the conformal interfaces with total transmission ($\lambda=1$), partial transmission/partial reflection ($\lambda=0.6$), and total reflection ($\lambda=1$). Here $\lambda$ is introduced in (6) and (7). We join the two decoupled chains at x=50x=-50, and the conformal interface is inserted at $x=0$. The two decoupled chains before joint together are defined on [400,50][-400,-50] and [50,400][-50,400], respectively.
Next, we will use (58) and (59) to extract $c_{\text{LR}}$ on a lattice model. The incident energy pulse in (57) will be generated by a local quantum quench. More concretely, we consider two decoupled free-fermion chains, one defined on sites from N-N to $N+N_{1}$, and the other defined on sites from $N+N_{1}+1$ to $N$.
The conformal interface is inserted locally at two sites 0 and $1$. That is, the Hamiltonians for the two decoupled chains are $H_{1}$ and $H_{2}$, where
@@
where $\{c_{i}^{\dagger},c_{j}\}=\delta_{ij}$ and $\{c_{i},c_{j}\}=\{c_{i}^{\dagger},c_{j}^{\dagger}\}=0$. $H_{2}$ has the same expression as (6), except that now the system is defined on sites from $N+N_{1}+1$ to $N$ and the conformal defect is defined on sites 0 and $1$.
@@
where $H_{\text{int}}=-\frac{1}{2}c_{N_{1}}^{\dagger}c_{N_{1}+1}+h.c.$. As shown in Fig.8, the excitations are generated from $x=-N+N_{1}$, where we join the two chains. Note that, here $H_{\rm int}$ matches the bulk hopping term on both the left and right chains and does not define an interface. The conformal defect is instead inserted in the right chain $H_{2}$. As time evolves, we have a right-moving pulse and a left-moving pulse. When the right-moving energy pulse arrives at the conformal interface (at $x=0$), it will split into two parts: one is transmitted and the other is reflected. The amount of energy transmitted/reflected depends on the details of the conformal interface, which is characterized by $\lambda$ in (7). As shown in Fig.8, when the conformal interface is totally transmissive ($\lambda=1$), the right-moving energy pulse propagates freely in space. If the conformal interface is partially transmissive ($0<\lambda<1$), the pulse is partially transmitted and partially reflected. If the interface is totally reflective ($\lambda=0$), then one can find that the energy pulse is totally reflected, which becomes left-moving.
@@ Appendix B Lattice calculation of $c_{\text{eff}}$
First, we consider two decoupled CFTs defined in [l,0][-l,0] and $[0,l]$ and join them at their ends at $x=0$ through a conformal interface. As shown in [54], the time evolution of entanglement entropy for the subsystem $A=[0,l]$ is described by
@@ Appendix C Liouvillian gap for general locations
Figure 11: The Liouvillian gap $g(\lambda)$ as a function of the total system size $L=2N$, with local dissipation applied at (a) $n_{d}=N-10$ and (b) $n_{d}=\frac{N}{2}$. For both the cases with ($\lambda=0.8$) and without ($\lambda=1$) conformal interface, we observe good cubic scaling, suggesting that it universally holds for general choices of the dissipative site $n_{d}$. The total system size is chosen as $2N=200,260,320,\dots,800$, and the dissipation strength is set to a small value of $\gamma=0.05$.
@@ Appendix D Dynamics of particle density
Figure 12: Time evolution of particle density $\langle c_{x}^{\dagger}c_{x}\rangle_{t}$ in the presence of local dissipation at a site $n_{d}$ in the left half chain. (a) The dissipation is applied at the left boundary $n_{d}=1$, and we examine the particle density at the same site $x=1$. (b) The dissipation is applied at at the left boundary $n_{d}=1$, and we examine the particle density at site $x=N+10=110$. (c) The dissipation is applied at the closest site to the interface $n_{d}=N-1$, and we examine the particle density at at the left boundary $x=1$. (d) The dissipation is applied at the closest site to the interface $n_{d}=N-1$, and we examine the particle density at site $x=N+10=110$. Here, the total system is of length $2N=200$ and the dissipation strength is set to $\gamma_{-}=0.1$. Here the initial states are set to be the ground state of the interface Hamiltonian, therefore the initial values of the particle density varies with the interface parameter $\lambda$.
@@
at site $x$ under the time evolution driven by the Lindbladian superoperator in Eq. (14) in the main text. The initial condition is chosen to be $\rho(t=0)=|\text{GS}\rangle\langle\text{GS}|$, where $|\text{GS}\rangle$ is the ground state of $H$ with the conformal interface as described in Eqs. (6) and  (7). A particle loss and/or gain is then introduced at site $n_{d}$. Here we include both losses and gains, with jump rates (dissipation strengths) $\gamma_{-}$ and $\gamma_{+}$, respectively [See (16)]. Since the Lindbladian superoperator is quadratic, the dynamics is exactly solvable, and the state remains Gaussian during the time evolution. In particular, the dynamics is fully described by the time-dependent correlation functions
@@
$G(t)=e^{\textbf{H}_{\text{eff}}t}G(0)e^{\textbf{H}_{\text{eff}}^{\dagger}t}+\int^{t}_{0}dze^{\textbf{H}_{\text{eff}}(t-z)}\Gamma^{+}e^{\textbf{H}_{\text{eff}}^{\dagger}(t-z)}.$ (73)
where $\textbf{H}_{\text{eff}}=i\textbf{H}-\frac{1}{2}(\mathbf{\Gamma}^{+}+\mathbf{\Gamma}^{-})$ is the (kernel of the) effective non-hermitian Hamiltonian as defined in Eq. (31) in the main text (but including both gain and loss here), and $\mathbf{\Gamma}^{\pm}_{x,y}=\gamma_{\pm}\delta_{x,y}\delta_{x,n_{d}}$ describes the particle gain and loss at the same site $n_{d}$.